OpenStudy (anonymous):

please help me Darrell has $55 to purchase a balloon bouquet for a celebration. A balloon bouquet costs $25 to make and $1.50 per balloon. How many balloons can Darrell have in the balloon bouquet? i think 30 flowere

OpenStudy (anonymous):

@peachpi @dan815

OpenStudy (anonymous):

25 + 1.50x = 55 -25 -25 1.50x = 30 x = 20

OpenStudy (anonymous):

@thatfoolkasey

OpenStudy (anonymous):

@thatfoolkasey

OpenStudy (anonymous):

y = 5

OpenStudy (anonymous):

because its a straight line across the y axis